# Controlling Ensemble Variance in Diffusion Models: An Application for Reanalyses Downscaling

### v1

@@ Abstract
In recent years, diffusion models have emerged as powerful tools for generating ensemble members in meteorology. In this work, we demonstrate that a Denoising Diffusion Implicit Model (DDIM) can effectively control ensemble variance by varying the number of diffusion steps. Introducing a theoretical framework, we relate diffusion steps to the variance expressed by the reverse diffusion process. Focusing on reanalysis downscaling, we propose an ensemble diffusion model for the full ERA5-to-CERRA domain, generating variance-calibrated ensemble members for wind speed at full spatial and temporal resolution. Our method aligns global mean variance with a reference ensemble dataset and ensures spatial variance is distributed in accordance with observed meteorological variability. Additionally, we address the lack of ensemble information in the CARRA dataset, showcasing the utility of our approach for efficient, high-resolution ensemble generation.
@@ 1 Introduction
The task is typically framed as an image-to-image prediction, transforming a low-resolution input into a high-resolution output. Widely adopted prediction-based models, such as U-Net [1] and ViT [2], operate deterministically and predict the most likely outcome for a given input. While this can produce high-quality predictions, these models lack the ability to quantify uncertainty or capture extreme events. To address this limitation, conditioned generative approaches, such as diffusion models [3], have gained traction in the scientific community. Unlike deterministic models, diffusion models sample outcomes from a learned distribution, producing diverse ensemble members even when conditioned on the same low-resolution input. These ensembles provide valuable insights into uncertainty estimation, the underlying physical system, and extreme events.
In this paper, we apply a Denoising Diffusion Implicit Model (DDIM) [4] to downscale meteorological reanalyses, focusing on wind speed. Specifically, we perform the transformation from a global reanalysis (ERA5 [5]) to a regional reanalysis (CERRA [6]) over the European domain. Building on prior work [7], we develop a diffusion model capable of operating across the full CERRA domain and we explore architectural modifications to scale diffusion models to large domains. A critical aspect of generative diffusion is the statistical properties of the generated distribution, particularly the variance. We aim to mathematically analyze how variance relates to the model’s architecture and ensure its spatial distribution aligns with physical expectations. Our analysis begins with a mathematical exploration of the reverse diffusion process, revealing that the variance of ensemble members relates to the number of diffusion steps. We experimentally validate this relationship in the ERA5-to-CERRA task, leveraging the CERRA-EDA [6] ensemble dataset for comparison. Our results demonstrate that adjusting the number of diffusion steps provides sufficient control to match the variance expressed by CERRA-EDA. This establishes our method as a practical tool for tuning ensemble diffusion to align with known uncertainty. Finally, studying the spatial distribution of the variance, we provide a methodology to obtain the optimal number of diffusion steps in specific applications, and we showcase our approach by applying it to a dataset of the Artic region lacking ensemble information, CARRA [8].
@@ 1.1 Background
Reanalysis datasets are a cornerstone of modern climate and weather research, providing gridded, high-resolution reconstructions of past atmospheric conditions. Reanalysis are widely used by researchers, institutions, and organizations to support studies ranging from climate variability to extreme event prediction [9, 10, 11]. High-resolution reanalyses are often created using downscaling techniques, which refine coarse global climate data into finer spatial and temporal scales. Traditional dynamical downscaling is computationally expensive, requiring significantly more resources than coarser global climate models and leading to data gaps and incomplete assessments of model uncertainty and regional climate variability [12, 13].
@@
In recent years, deep learning has emerged as an effective alternative for downscaling, offering powerful and cost-effective methods for producing high-resolution data [14, 15]. A variety of deep learning models for downscaling have been proposed [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], showcasing their ability to capture complex spatial patterns and generate high-quality outputs across a range of meteorological variables.
@@
Among these methods, diffusion models stand out for their ability to capture the uncertainty of the downscaling process, enabling not only the generation of high-resolution data but also produce probabilistic ensembles by sampling from a latent space. This capability makes them highly effective for modeling uncertainty and variability, which are crucial in meteorological applications [29, 30, 31, 32, 33, 34, 35].
@@
Downscaled ensemble datasets are essential for quantifying uncertainties [36, 37, 38], detecting extreme events [39], and performing ensemble-based statistical analyses critical for assessing the likelihood of compound extreme events [40, 41]. In recent years an increasing amount of publications focused on the utilization of generative deep learning architectures in the downscaling task, utilizing the inherently probabilistic nature of the models to generate ensemble members.
In [20], the authors extend generative adversarial networks (GANs) to downscale precipitation forecasts, addressing the challenges of adding fine-scale structure and correcting forecast errors. Using high-resolution radar data as ground truth, the study demonstrates that GANs and VAE-GANs can produce spatially coherent precipitation maps while maintaining statistical properties comparable to state-of-the-art post-processing methods. In [32], the authors propose using deep generative diffusion models to emulate and enhance probabilistic weather forecasting. Learning from historical data, diffusion model generate realistic ensemble forecasts at a fraction of the computational cost of traditional physics-based simulations. The generated ensembles replicate the statistical properties and predictive skill of operational forecasts, while also correcting biases, improving reliability, and better capturing probabilities of extreme weather events. This scalable approach has potential applications in both weather forecasting and large-scale climate risk assessment. In [18], the authors propose a statistical downscaling method for satellite sea surface wind (SSW) data using generative adversarial networks (GANs) with a dual learning scheme. The dual learning approach combines high-resolution SSW reconstruction with degradation kernel estimation in a closed loop. In [38], the authors propose a cascading deep learning framework based on generative diffusion models for downscaling tropical cyclone (TC) wind fields from low-resolution ERA5 data (0.25°) to high-resolution observations (0.05°) at 6-hour intervals. The framework consists of a debiasing neural network to improve wind speed accuracy and a conditional denoising diffusion probabilistic model (DDPM) for super-resolution and ensemble generation. Applied to 85 TC events, the model achieves realistic spatial structures, enabling accurate high-resolution TC wind modeling and risk assessment. In [37], the authors present a hybrid downscaling framework combining dynamical downscaling with generative diffusion models to enhance resolution and improve uncertainty estimates for regional climate projections. The method uses a regional climate model (RCM) to downscale Earth System Model (ESM) outputs to intermediate resolution, followed by a diffusion model for further refinement. Applied to the CMIP6 ensemble, this approach provides more accurate uncertainty bounds, lower errors than bias correction and spatial disaggregation (BCSD), and better captures spectra and multivariate correlations. The framework offers a cost-effective and accurate alternative for downscaling large climate projection ensembles.
@@ 1.2 Diffusion models basics
Diffusion models, and in particular Denoising Diffusion Implicit Models (DDIM) [4], are state-of-the-art generative models able to generate highly realistic data from noise. In particular, given a sequence $\{\alpha_{t}\}_{t=1}^{T}\subseteq(0,1)$ such that $\alpha_{0}\approx 0$ and $\alpha_{T}\approx 1$, called noise schedule, a DDIM defines an iterative procedure to turn random noise $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ into a new sample $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$, where $p_{gt}(\boldsymbol{x}_{T})$ represents the probability distribution from which the training data is sampled [^1] Note that here we use a slightly different notation compared to e.g. [4], as we consider the time running reversely. We made this choice as this notation simplifies the mathematical derivation of our formula in section 3.1.. This procedure is obtained through a simple yet effective idea: let $\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T-1},\boldsymbol{x}_{T}\}$ be a sequence such that $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$, and $\boldsymbol{x}_{t}\sim p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{T})$ for any $t=T-1,\dots,0$. If $p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{T})=\mathcal{N}(\sqrt{\alpha_{t}}\boldsymbol{x}_{T},(1-\alpha_{t})\boldsymbol{I})$, then:
@@
After $T$ steps, this process leads to a new sample $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$. For a detailed introduction to DDIM, please refer to [4, 42, 43].
@@ 2 Data
In this section we describe the primary datasets used in this study. Our task involves using the global reanalyses ERA5 [44] as the low-resolution input to condition our model in predicting a high-resolution downscaled counterpart, trained on the regional reanalyses CERRA [6]. The resulting ensemble members are validated against CERRA-EDA, an ensemble version of CERRA at half the temporal and spatial resolution. We will also apply our methods to CARRA, a dataset built on the same core model as CERRA and that covers the Arctic region, but lack ensemble members. This section begins with a brief introduction to ERA5, followed by a focused discussion on CERRA, CARRA, and CERRA-EDA.
@@ 2.1 ERA5
ERA5 [5], the fifth generation ECMWF reanalysis, is a globally recognized dataset extensively utilized in climate and atmospheric research. Spanning from 1940 to the present, it provides hourly estimates of atmospheric, land, and oceanic variables at a 0.25° horizontal resolution (approximately 30 km) and 137 vertical levels. Based on the ECMWF Integrated Forecasting System, ERA5 incorporates advanced numerical weather prediction models and satellite data, ensuring high reliability and precision [44]. Its widespread use spans diverse applications, including precipitation trends, temperature analysis, wind studies, and extreme event research, making it an indispensable resource in climate science.
@@ 2.2 CERRA
The Copernicus Regional Reanalysis for Europe (CERRA [6]) is a sophisticated high-resolution Regional ReAnalysis (RRA) dataset specifically designed for the European region. It is a product of the European Copernicus Climate Change Service (C3S), executed through a contract with the Swedish Meteorological and Hydrological Institute (SMHI), in collaboration with subcontractors Meteo-France and the Norwegian Meteorological Institute. CERRA operates at a high 5.5 km horizontal resolution, offering enhanced spatial detail for meteorological variables across Europe.
@@ 2.3 CARRA
The C3S Arctic Regional Reanalysis (CARRA) dataset [8] addresses the unique environmental and climatic conditions of the Arctic region, which are insufficiently captured by global reanalyses such as ERA5. While ERA5 provides a 31 km global resolution, CARRA offers 2.5 km horizontal resolution, capturing finer details essential for understanding the Arctic’s rapidly changing climate. Observational records and climate projections show that warming in the Arctic occurs at over twice the global average rate, leading to increased environmental and economic activities and driving a need for detailed data to support climate adaptation and management in the region.
@@ 2.4 CERRA-EDA
CERRA-EDA [45] is an Ensemble of Data Assimilation (EDA) system comprising 10 members. It is based on the same code as the deterministic CERRA system and uses the same observational data, with the exception of the Microwave Sounding Unit and has a lower 11-km horizontal resolution.
@@ 3.1 Signal and noise rates
Figure 2: Reverse diffusion signal and noise ratios with cosine scheduling, including the reduced signal rate
In Section 1 we discussed the importance of the diffusion schedule $\{\alpha_{t}\}_{t=1,\dots,T}$ in the generation process of DDIM. Recently, various schedules have been proposed to find a perfect balance between the signal and noise in the diffusion process, including linear schedules [3] and sinusoidal schedules [46, 4]. In this study, we adopt sinusoidal functions due to their ability to achieve smooth and controlled transitions between signal and noise components, which is crucial for effective denoising and high-quality image generation. In particular, we consider the following choice for the signal rate as in [42]:
@@
for which it holds that $\left(\sqrt{\alpha_{t}}\right)^{2}+\left(\sqrt{1-\alpha_{t}}\right)^{2}=\sin\left(\pi\frac{T-t}{T}\right)^{2}+\cos\left(\pi\frac{T-t}{T}\right)^{2}=1$, ensuring that the noisy images, formed as a weighted sum of signal and noise, consistently maintain unit variance [46, 47]. The extremes of the schedule (i.e. when $t=0$ and $t=T$) can be problematic, as near-pure signal or noise levels often cause instability. Following common practice [47, 48], we clamp the extremes to maintain a balanced Signal-to-Noise Ratio (SNR), preventing degenerate cases and ensuring stable training and reliable reverse diffusion.

### v2

@@ Abstract
In recent years, diffusion models have emerged as powerful tools for generating ensemble members in meteorology. In this work, we demonstrate how a Denoising Diffusion Implicit Model (DDIM) can effectively control ensemble variance by varying the number of diffusion steps. Introducing a theoretical framework, we relate diffusion steps to the variance expressed by the reverse diffusion process. Focusing on reanalysis downscaling, we propose an ensemble diffusion model for the full ERA5-to-CERRA domain, generating variance-calibrated ensemble members for wind speed at full spatial and temporal resolution. Our method aligns global mean variance with a reference ensemble dataset and ensures spatial variance is distributed in accordance with observed meteorological variability. Additionally, we address the lack of ensemble information in the CARRA dataset, showcasing the utility of our approach for efficient, high-resolution ensemble generation.
@@ 1 Introduction
The task is typically framed as an image-to-image prediction, transforming a low-resolution input into a high-resolution output. Widely adopted prediction-based models, such as U-Net [ronneberger2015] and ViT [liu2021swin], operate deterministically and predict the most likely outcome for a given input. While this can produce high-quality predictions, these models lack the ability to quantify uncertainty or capture extreme events. To address this limitation, conditioned generative approaches, such as diffusion models [ho2020denoising], have gained traction in the scientific community. Unlike deterministic models, diffusion models sample outcomes from a learned distribution, producing diverse ensemble members even when conditioned on the same low-resolution input. These ensembles provide valuable insights into uncertainty estimation, the underlying physical system, and extreme events.
In this paper, we apply a Denoising Diffusion Implicit Model (DDIM) [song2022denoisingdiffusionimplicitmodels] to downscale meteorological reanalyses, focusing on wind speed. Specifically, we perform the transformation from a global reanalysis (ERA5 [Hersbach_ERA5_2018]) to a regional reanalysis (CERRA [ridal2024cerra]) over the European domain. Building on prior work [Merizzi2024], we develop a diffusion model capable of operating across the full CERRA domain and we explore architectural modifications to scale diffusion models to large domains. A critical aspect of generative diffusion is the statistical properties of the generated distribution, particularly the variance. We aim to mathematically analyze how variance relates to the model’s architecture and ensure its spatial distribution aligns with physical expectations. Our analysis begins with a mathematical exploration of the reverse diffusion process, revealing that the variance of ensemble members relates to the number of diffusion steps. Although previous works have already reported a reduction in variability when using fewer diffusion steps [song2022denoisingdiffusionimplicitmodels, nichol2021improved] the best of our knowledge, no study has systematically investigated the relationship between ensemble variance and the number of diffusion steps. We experimentally validate this relationship in the ERA5-to-CERRA task, leveraging the CERRA-EDA [ridal2024cerra] ensemble dataset for comparison. Our results demonstrate that adjusting the number of diffusion steps provides sufficient control to match the variance expressed by CERRA-EDA. This establishes our method as a practical tool for tuning ensemble diffusion to align with known uncertainty. Finally, studying the spatial distribution of the variance, we provide a methodology to obtain the optimal number of diffusion steps in specific applications, and we showcase our approach by applying it to a dataset of the Artic region lacking ensemble information, CARRA [Copernicus_Arctic_Reanalysis].
@@ 1.1 Background
Reanalysis datasets are a cornerstone of modern climate and weather research, providing gridded, high-resolution reconstructions of past atmospheric conditions. Reanalysis are widely used by researchers, institutions, and organizations to support studies ranging from climate variability to extreme event prediction [rs12060980, MOREIRA2019335, zandler2020reanalysis]. High-resolution reanalyses are often created using downscaling techniques, which refine coarse global climate data into finer spatial and temporal scales. Traditional dynamical downscaling is computationally expensive, requiring significantly more resources than coarser global climate models and leading to data gaps and incomplete assessments of model uncertainty and regional climate variability [Pierce2009, Goldenson2023].
@@
In recent years, deep learning has emerged as an effective alternative for downscaling, offering powerful and cost-effective methods for producing high-resolution data [Sun2024, perez2024]. A variety of deep learning models for downscaling have been proposed [vandal2017deepsd, cheng2022, rs14030769, Ji2020Downscaling, Harris2022, passarella2022reconstructing, quesada2022repeatable, ZHANG2021119321, nguyen2023climatelearn, Harder2022Generating, dujardin2022wind, 9932682, price2023gencast], showcasing their ability to capture complex spatial patterns and generate high-quality outputs across a range of meteorological variables.
@@
Among these methods, diffusion models stand out for their ability to capture the uncertainty of the downscaling process, enabling not only the generation of high-resolution data but also produce probabilistic ensembles by sampling from a latent space. This capability makes them highly effective for modeling uncertainty and variability, which are crucial in meteorological applications [mardani2024, lopezgomez2024, ling2024diffusion, Li2024, wan2023debias, asperti2023precipitation, leinonen2023latent].
On the other hand, despite their early success in generative tasks, Generative Adversarial Networks (GANs) have proven to be highly unstable and difficult to generalize [saad2024survey, wiatrak2019stabilizing]. Their adversarial training often leads to mode collapse, where the generator produces limited variability, and convergence is not theoretically guaranteed [cobbinah2025diversity]. Furthermore, GAN-based approaches tend to be problem-specific, requiring extensive architectural and hyperparameter tuning to maintain stability [saxena2021generative], and numerous variants have been proposed merely to address training issues. As a result, GANs are being increasingly phased out in climate and meteorological downscaling tasks [rampal2024enhancing, dhariwal2021diffusion] and are being superseded by diffusion models, which offer stable likelihood-based training, controllable stochasticity, and consistent convergence behavior.
@@
Downscaled ensemble datasets are essential for quantifying uncertainties [Wootten2020The, lopez2024dynamical, Lockwood2024], detecting extreme events [trenberth2015attribution], and performing ensemble-based statistical analyses critical for assessing the likelihood of compound extreme events [Kornhuber2023, Anderson2019]. In recent years an increasing amount of publications focused on the utilization of generative deep learning architectures in the downscaling task, utilizing the inherently probabilistic nature of the models to generate ensemble members.
Recent studies have explored the use of generative models for meteorological downscaling and ensemble generation. In [Harris2022], generative adversarial networks (GANs) were extended to downscale precipitation forecasts, producing spatially coherent maps while correcting forecast errors. Similarly, [rs14030769] applied a dual-learning GAN framework to satellite sea surface wind (SSW) data, combining high-resolution reconstruction with degradation kernel estimation. While these GAN-based approaches demonstrated promising results, they remained limited in their ability to model uncertainty.
More recently, diffusion-based generative models have gained attention for probabilistic weather forecasting and downscaling. [Li2024] employed diffusion models to generate ensemble forecasts that replicate the statistical properties and predictive skill of operational systems, while improving bias correction and reliability. In [Lockwood2024], a conditional diffusion model was proposed for tropical cyclone wind-field downscaling, achieving realistic high-resolution structures and ensemble variability. Finally, [lopez2024dynamical] combined dynamical and diffusion-based downscaling to enhance regional climate projections, improving uncertainty estimates and multivariate correlations.
Together, these works highlight the growing relevance of diffusion models for meteorological ensemble generation. However, none of them explicitly investigate how the variance of generated ensembles can be controlled or calibrated—a gap that the present work addresses through a theoretical and empirical analysis of the relationship between variance and the number of diffusion steps.
@@ 1.2 Diffusion models basics
Diffusion models, and in particular Denoising Diffusion Implicit Models (DDIM) [song2022denoisingdiffusionimplicitmodels], are state-of-the-art generative models able to generate highly realistic data from noise. In particular, given a sequence $\{\alpha_{t}\}_{t=1}^{T}\subseteq(0,1)$ such that $\alpha_{0}\approx 0$ and $\alpha_{T}\approx 1$, called noise schedule, a DDIM defines an iterative procedure to turn random noise $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ into a new sample $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$, where $p_{gt}(\boldsymbol{x}_{T})$ represents the probability distribution from which the training data is sampled [^1] Note that here we use a slightly different notation compared to e.g. [song2022denoisingdiffusionimplicitmodels], as we consider the time running reversely. We made this choice as this notation simplifies the mathematical derivation of our formula in section 3.1.. This procedure is obtained through a simple yet effective idea: let $\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T-1},\boldsymbol{x}_{T}\}$ be a sequence such that $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$, and $\boldsymbol{x}_{t}\sim p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{T})$ for any $t=T-1,\dots,0$. If $p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{T})=\mathcal{N}(\sqrt{\alpha_{t}}\boldsymbol{x}_{T},(1-\alpha_{t})\boldsymbol{I})$, then:
@@
After $T$ steps, this process leads to a new sample $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$. For a detailed introduction to DDIM, please refer to [song2022denoisingdiffusionimplicitmodels, asperti2023image, karras2022elucidating].
@@ 2 Data
In this section we describe the primary datasets used in this study. Our task involves using the global reanalyses ERA5 [hersbach2020era5] as the low-resolution input to condition our model in predicting a high-resolution downscaled counterpart, trained on the regional reanalyses CERRA [ridal2024cerra]. The resulting ensemble members are validated against CERRA-EDA, an ensemble version of CERRA at half the temporal and spatial resolution. We will also apply our methods to CARRA, a dataset built on the same core model as CERRA and that covers the Arctic region, but lack ensemble members. The respective domains of CERRA and CARRA are reported in Figure 1. This section begins with a brief introduction to ERA5, followed by a focused discussion on CERRA, CARRA, and CERRA-EDA.
@@ 2.1 ERA5
ERA5 [Hersbach_ERA5_2018], the fifth generation ECMWF reanalysis, is a globally recognized dataset extensively utilized in climate and atmospheric research. Spanning from 1940 to the present, it provides hourly estimates of atmospheric, land, and oceanic variables at a 0.25° horizontal resolution (approximately 30 km) and 137 vertical levels. Based on the ECMWF Integrated Forecasting System, ERA5 incorporates advanced numerical weather prediction models and satellite data, ensuring high reliability and precision [hersbach2020era5]. Its widespread use spans diverse applications, including precipitation trends, temperature analysis, wind studies, and extreme event research, making it an indispensable resource in climate science.
@@ 2.2 CERRA
The Copernicus Regional Reanalysis for Europe (CERRA [ridal2024cerra]) is a sophisticated high-resolution Regional ReAnalysis (RRA) dataset specifically designed for the European region. It is a product of the European Copernicus Climate Change Service (C3S), executed through a contract with the Swedish Meteorological and Hydrological Institute (SMHI), in collaboration with subcontractors Meteo-France and the Norwegian Meteorological Institute. CERRA operates at a high 5.5 km horizontal resolution, offering enhanced spatial detail for meteorological variables across Europe.
@@ 2.3 CARRA
The C3S Arctic Regional Reanalysis (CARRA) dataset [Copernicus_Arctic_Reanalysis] addresses the unique environmental and climatic conditions of the Arctic region, which are insufficiently captured by global reanalyses such as ERA5. While ERA5 provides a 31 km global resolution, CARRA offers 2.5 km horizontal resolution, capturing finer details essential for understanding the Arctic’s rapidly changing climate. Observational records and climate projections show that warming in the Arctic occurs at over twice the global average rate, leading to increased environmental and economic activities and driving a need for detailed data to support climate adaptation and management in the region.
@@ 2.4 CERRA-EDA
CERRA-EDA [Ridal2024] is an Ensemble of Data Assimilation (EDA) system comprising 10 members. It is based on the same code as the deterministic CERRA system and uses the same observational data, with the exception of the Microwave Sounding Unit and has a lower 11-km horizontal resolution.
@@ 3.1 Signal and noise rates
Figure 2: Reverse diffusion signal and noise ratios with cosine scheduling, including the reduced signal rate.
In Section 1 we discussed the importance of the diffusion schedule $\{\alpha_{t}\}_{t=1,\dots,T}$ in the generation process of DDIM. Recently, various schedules have been proposed to find a perfect balance between the signal and noise in the diffusion process, including linear schedules [ho2020denoising] and sinusoidal schedules [nichol2021improved, song2022denoisingdiffusionimplicitmodels]. In this study, we adopt sinusoidal functions due to their ability to achieve smooth and controlled transitions between signal and noise components, which is crucial for effective denoising and high-quality image generation. In particular, we consider the following choice for the signal rate as in [asperti2023image]:
@@
for which it holds that $\left(\sqrt{\alpha_{t}}\right)^{2}+\left(\sqrt{1-\alpha_{t}}\right)^{2}=\sin\left(\pi\frac{T-t}{T}\right)^{2}+\cos\left(\pi\frac{T-t}{T}\right)^{2}=1$, ensuring that the noisy images, formed as a weighted sum of signal and noise, consistently maintain unit variance [nichol2021improved, Chen2023]. The extremes of the schedule (i.e. when $t=0$ and $t=T$) can be problematic, as near-pure signal or noise levels often cause instability. Following common practice [Chen2023, Hoogeboom2023], we clamp the extremes to maintain a balanced Signal-to-Noise Ratio (SNR), preventing degenerate cases and ensuring stable training and reliable reverse diffusion.